Please I want some to do answer it clearly and step by step

3. A potter works with a uniform solid grindstone wheel that has a mass of 18 kg and a radius of 25 cm. The grindstone initially rotates at 1200 revlmin. The grindstones motor is turned off and the potter presses an instrument perpendicular against the grindstone with a force of 40 N to bring it to a stop. The coefficient of friction with the grindstone is 0.7. The grindstone can be considered a uniform solid cylinder for which I mr (a) (10 points) What is the torque on the grindstone in stopping it? (b) (10 points) Find the angular deceleration of the grindstone. (c) (10 points) How long does it take the grindstone to stop?
Answer #1

a)

frictional force

f =uF =0.7*40 =28 N

Torque acting on the grindstone to stop it

T=fR =28*0.25 =7 N-m

b)

Moment of inertia

I=(1/2)*18*0.252=0.5625 Kg-m2

SInce

T=I*alpha

angular deceleration is

alpha=T/I =7/0.5625 =12.44 rad/sec2

c)

from

W=Wo+alpha*t

since grinding wheel stops ,therefore W=0 ,since it is decelerating alpha is negative

0=1200*(2pi/60)-12.44*t

t=10.1 sec